Circular No. 6458

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 Q1

The following CCD precise positions have been reported:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 21.43560 4^{h}47^{m}16.^{s}21 −8^{o}49′08′′.6 14.6 Offutt
21.47828 4 47 20.20 −8 48 16.2 14.2  _{”}
21.77022 4 47 48.26 −8 42 12.6  Kojima
21.77206 4 47 48.35 −8 42 09.4  Abe
21.77323 4 47 48.48 −8 42 09.0  Kojima
21.80573 4 47 51.69 −8 41 29.3  Abe
22.12717 4 48 22.46 −8 34 44.2  Pravec
22.12979 4 48 22.74 −8 34 41.1  _{”}

W. Offutt (Cloudcroft). 0.60-m f/7 Ritchey-Chrétien. T. Kojima (Chiyoda). 0.25-m f/6 reflector; some interfering clouds. H. Abe (Yatsuka). 0.26-m f/6.0 reflector. P. Pravec (Ondřejov). 0.65-m f/3.6 reflector.

COMET C/1995 O1 (HALE-BOPP) N. Biver, Observatoire de Paris-Meudon (OPM); A. Enzian, Labora-

toire de Glaciologie, Grenoble; D. Bockelée-Morvan, P. Colom, J. Crovisier, E. Gérard, E. Lellouch, and H. Rauer, OPM; D. Despois, Observatoire de Bordeaux; R. Moreno and G. Paubert, Institut de Radio Astronomie Millimétrique (IRAM), Granada; and J. K. Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo, report the detection of methyl cyanide (CH_{3}CN) in comet C/1995 O1 with the IRAM 30-m telescope: “We detected on Aug. 14–17 the 8(0)-7(0), 8(1)-7(1), 8(2)-7(2), and 8(3)-7(3) lines of CH_{3}CN around 147 GHz with integrated line areas on a Tb scale of 0.071 ± 0.008, 0.073 ± 0.008, 0.038 ± 0.008, and 0.051 ± 0.009 K km/s, respectively. We derive a CH_{3}CN production rate of 1.2 × 10^{26} molecules/s. For comparison, Q[CO], Q[HCN], and Q[OH] were 1.5 × 10^{29}, 9 × 10^{26}, and 2.5 × 10^{29} molecules/s, respectively, as measured at the same time at IRAM and Nançay.”

SUPERNOVA 1996aq IN NGC 5584 Corrigendum. On IAUC 6454, line 19, for FWZI read HWZI

1996 August 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green